OpenStudy (anonymous):

Determine whether u and v are orthogonal, parallel, or neither.u = <15,45> v= <-5,12>

OpenStudy (anonymous):

Is it neither?

OpenStudy (anonymous):

Do you know any tests for orthogonality?

OpenStudy (anonymous):

I know it's not orthogonal

OpenStudy (anonymous):

It doesn't equal zero when you multiply them out

OpenStudy (anonymous):

alright, what about parallel? how can you tell when two vectors are pointing in the same direction?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

To test if they're parallel, I have to see if Cos in the equation equals to 1 ot -1

OpenStudy (anonymous):

*or

OpenStudy (anonymous):

That works too. If cos(theta) = 1 or -1, that means the angle between the two vectors is either 0 or pi (180), and the vectors are pointing in the same direction, or in opposite directions (negative direction)

OpenStudy (anonymous):

When I used the equaltion: u ▪ v = ||u|| ||v||cos I didn't get 1 or -1. I need to know if I'm right

OpenStudy (anonymous):

If you didnt get 1 or -1, then you already know they arent parallel, so it must be neither. They are skew vectors.

OpenStudy (anonymous):

I need someone to check my work xD just in case

OpenStudy (anonymous):

its right. those vectors are pointing in totally different directions. There's no way they are parallel.

OpenStudy (anonymous):

can you show your work to prove it?

OpenStudy (anonymous):

lets see.... \[||u||=13,||v||=15\sqrt{10}\]\[u\cdot v=-75+540=465\]\[\frac{u\cdot v}{||u||\cdot ||v||}\ne \pm 1\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

And then you would have 465/ 517.45

OpenStudy (anonymous):

but you can draw the two vectors and see they arent parallel.

OpenStudy (anonymous):

then you set it up like: Cos theta = 465/ 517.45 =26 degress?

OpenStudy (anonymous):

*Degrees. Anyway, that's why I'm confused?

OpenStudy (anonymous):

Shouldn't it = 1 or -1 when you take the cos of it?

OpenStudy (anonymous):

as soon as you show that cos(theta) isnt 1 or -1, you are done. The question didnt ask for the angle between the two. You showed they arent perpendicular since their dot product isnt 0. You showed they arent parallel since cos(theta) doesnt equal 1 or -1. So it must be neither. There is nothing more you need to show.

OpenStudy (anonymous):

SO...I WAS RIGHT :D

OpenStudy (anonymous):

Thank you for reassuring me